### v1

# Few-Shot Learner Generalizes Across AI-Generated Image Detection

Abstract
Current fake image detectors trained on large synthetic image datasets perform satisfactorily on limited studied generative models. However, they suffer a notable performance decline over unseen models. Besides, collecting adequate training data from online generative models is often expensive or infeasible. To overcome these issues, we propose Few-Shot Detector (FSD), a novel AI-generated image detector which learns a specialized metric space to effectively distinguish unseen fake images by utilizing very few samples. Experiments show FSD achieves state-of-the-art performance by $+7.4\%$ average ACC on GenImage dataset. More importantly, our method is better capable of capturing the intra-category common features in unseen images without further training.
Deepfake Detection, Synthetic Image Detection, Few-Shot Learning
1 Introduction
The development of generative models has led to significant strides in synthesizing photorealistic images, making it much difficult to distinguish AI-generated images from real ones. Over the past few years, diffusion-based models (Ho et al., 2020) have exceeded GANs (Goodfellow et al., 2014) and become the main stream paradigm of image generation, due to their strong ability of synthesizing images of high-quality. With the rapid release of open-source models and web APIs, users can easily create unrealistic images with brief texts. However, the increasing abuse of deepfake technology has dramatically influenced the human society and sparked unprecedented concerns on the credibility of online information.
To curb the spread of AI-generated images used for malicious purposes, there is a growing urgency to develop detection methods for distinguishing fake images from authentic ones. Early works (Zhang et al., 2019b; Chai et al., 2020) focus on identifying artifacts in images, which prove to be particularly useful for detecting GAN-generated images. With the improvement of synthetic image datasets (Zhu et al., 2023; Boychev & Cholakov, 2024), detecting fake images from seen models is no longer difficult. Thus, a new challenge is to develop a classifier that generalizes across diverse unseen generative models. To address this problem, some studies (Ojha et al., 2023; Liu et al., 2024) adopt a large vision-language model to analyze images from a semantic perspective, while others (Wang et al., 2023; Luo et al., 2024) leverage the reconstruction ability of diffusion models.
Another difficulty is the acquisition of training data. It is often expensive or infeasible to collect enough images from closed-source models such as DALL-E (Ramesh et al., 2021) and Midjourney (mid, 2022), due to their service prices and access restrictions. However, most existing fake image detectors (Wang et al., 2023; Ojha et al., 2023) require large amounts of training data to achieve good generalization, while fine-tuning without adequate data often results in overfitting. Furthermore, the ongoing release and updating of generative models pose a significant challenge for any trained detectors to keep up.
After gaining a deep insight into these limitations, we propose a distinctive scheme based on few-shot learning to overcome them and design a comprehensive model to validate the efficacy of our solution. We first point out that current AI-generated image detection is a domain generalization task. Previous studies have been dedicated to finding a common indicator effective for all generative images. ̵However, they overlook the significant distinctions among data from different domains. We observe that images from unseen domains can actually be obtained in many real-world scenarios. Based on this fact, the complicated task can be transformed to a slightly easier one, called few-shot classification, by using relative few samples from unseen domains. Thus, we can extract rich domain information from these samples and use it to generalize across the unseen data.
Based on the above analysis, we propose Few-Shot Detector (FSD), a novel method which can detect fake images with few given samples. In the few-shot classification task, synthetic images from different generative models are classified into separate categories, while the real images are classified into one class. The goal is to train a classifier capable of generalizing to unseen classes not in the training set, with only a small number of examples from them. To solve this problem, FSD utilizes a prototypical network (Snell et al., 2017), which learns a metric space to be able to handle unseen data. Different from end-to-end classification methods, FSD requires receiving a few samples from several classes to compute the corresponding prototypical representations. The test image is then classified with the nearest-neighbor method. Thus, the bias in unseen classes which can not be learned from the training set can now be automatically applied to the classifier without extra training.
Experiments show that FSD obtains notable improvements over the state-of-the-art method by $+7.4\%$ average accuracy on GenImage dataset. We observe that only $10$ samples from the unseen classes can significantly improve the detection performance, which shows great privilege of our method. We also show that our approach is better equipped to keep pace with the rapid advancements in image generation, thereby highlighting the effectiveness of categorizing generated images by their sources. Our main contributions are as follows:
  • To the best of our knowledge, we are the first to reconceptualize AI-generated image detection as a few-shot classification task, bringing it closer to the real-world scenario applications.
  • We introduce an innovative synthetic image detection method, named Few-Shot Detector (FSD). Our approach is able to utilize few samples from unseen domains to achieve better detection performance.
  • Our experiments demonstrate the generalization ability of FSD, which significantly outperforms current state-of-the-art methods and can easily deal with new generative models without further training.
2 Related Work
2.1 Image Synthesis
Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) have revolutionized the field of generative modeling over the last decade. By simultaneously training a generator and a discriminator in an adversarial manner, GANs can generate samples of high resolution and good quality (Karras et al., 2020; Brock et al., 2019). However, they are prone to mode collapse, making it difficult for them to capture the full data distribution. An alternative technical route is likelihood-based modeling, which estimates the parameters of data distribution by maximizing the likelihood function. Within this framework, variational autoencoders (VAEs) (Kingma & Welling, 2014) ensure a smooth latent space which facilitates interpolation and exploration of new samples. Meanwhile, normalizing flow models (Papamakarios et al., 2021) allow for exact computation of the log-likelihood, which enables them to capture more complex data distributions. However, neither of them can match the high-quality image generation capabilities of GANs.
Diffusion Models (Ho et al., 2020; Song et al., 2021) have recently achieved remarkable performances in image synthesis. By reversing a gradual noising process, diffusion models can generate an image from random noise with several denoising steps. Ho & Salimans (2022) introduce a classifier-free guidance that enables the generation of images from text descriptions, eliminating the need for training a separate classifier. Another significant advancement is latent diffusion model (LDM) (Rombach et al., 2022), which applies diffusion process in latent space with a powerful autoencoder. This approach greatly enhances the visual fidelity of synthetic images while substantially reducing the required computational resources. Based on vast amounts of training data and substantial computing resources, large latent diffusion models (Podell et al., 2024; Esser et al., 2024) have become the most potent technique in image generation and the main source of synthetic images at present.
Figure 1: Detection pipeline of Few Shot Detector. FSD first maps the real and fake images from support set to the metric space by prototypical network and calculates the representation of each type. A test image is then mapped into the same space and classified according to the principle of nearest neighbor.
2.2 AI-generated Image Detection
Model-agnostic detection method. Learning based methods have been applied to detecting fake images from GANs (Zhang et al., 2019b; Qian et al., 2020; Chai et al., 2020). They treat synthetic image detection as a binary classification task, aiming to find the synthetic clues in generated images. Many studies (Zhang et al., 2019a; Gragnaniello et al., 2021) have found that classifiers do not generalize over unseen generative models, bringing generalization ability to the primary research objective. Meanwhile, Wang et al. (2020) reveal that detectors for CNN-generated images can be surprisingly generalizable due to the common systematic flaws shared among CNNs. Based on this work, Liu et al. (2023) analyze multi-view features to learn a robust representation, while Tan et al. (2024b) focus on high-frequency information to learn the source-agnostic features.
However, these common cues disappear when it comes to diffusion models, leading to the poor performance of earlier methods when applied to them. Therefore, more attention has been paid to semantic level. For instance, Ojha et al. (2023) prove that the informative image features from a pretrained CLIP model (Radford et al., 2021) can successfully be used for detecting fake images. Other studies (Liu et al., 2024; Tan et al., 2024a) deeply integrate textual information with images, thereby further enhancing the generalization capability. Moreover, patch-based detection method (Chen et al., 2024) has been proposed to reduce computational resources and accelerate detection speed. This method extracts features from a few patches instead of the whole image, enabling more efficient discrimination between fake and real images.
Diffusion-based detection method. The capability of open-source diffusion models to reconstruct images has opened a new technological route. DIRE (Wang et al., 2023) first uses the reconstruction loss to train a classifier, based on the observation that synthetic images can be reconstructed better than real ones. Following this discovery, SeDID (Ma et al., 2023) analyzes the loss errors between noised and denoised features at specific steps of processing, while LARE2 (Luo et al., 2024) focuses on the single-step reconstruction result in latent space. Both of them are able to improve the detection speed by not fully completing the entire reconstruction process. FakeInversion (Cazenavette et al., 2024) uses a BLIP (Li et al., 2022) to obtain a text description of the image for better reconstruction, which is capable of detecting images from unseen high-fidelity generators. AEROBLADE (Ricker et al., 2024) evaluates the autoencoder (AE) in different latent diffusion models and presents a training-free method for detecting fake images by AE reconstruction errors. These approaches achieve great success on open-source generative models. However, the differences between various diffusion models still can not be ignored, which makes it particularly challenging for them to detect synthetic images produced by numerous generative models, especially those that are closed-source.
3 Method
(a) training
(b) testing
Figure 2: Prototypical network process. During training, the optimal center of each class is learned by minimizing the distances between query samples and their corresponding centers. During testing, The test image is only classified between real and fake.
In this section, we first describe the few-shot classification task of synthetic image detection. Then we elaborate on our novel representation, namely Few-Shot Detector (FSD), as illustrated in Fig. 1.
3.1 Few-Shot Synthetic Image Detection
Few-shot classification is a specialized machine learning task, in which the classifier is challenged to generalize to entirely new classes that are not present in the training set, using only a limited number of examples from these novel classes. Different from previous studies that categorize all the synthetic images produced by different generative models into a single fake class, we classify the images based on their generators. That is, we treat each group of them as a distinct class. In contrast, all the authentic images are classified into a single class. During detection, the set of additional labeled images used to support the classification task is known as the support set, while the set of those to be tested is referred to as the query set. The goal of this task is to determine which category each image in the query set belongs to, based on the support set.
In detail, we are given a support set comprising samples from $N$ new classes $\mathcal{S}=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\cdots\cup\mathcal{S}_{N}$, where $\mathcal{S}_{i}=\{(\mathbf{x}_{i1},y_{i1}),(\mathbf{x}_{i2},y_{i2}),\ldots,(\mathbf{x}_{iK},y_{iK})\}$ is the image subset of class $i$. $K$ is the number of samples in each class. $\mathbf{x}_{ij}\in\mathbb{R}^{D}$ is an image feature of $D$ dimension and $y_{ij}=i$ is the corresponding label. Each unlabeled image in the query set will be assigned to one of the $N$ classes. And this is the so-called $N$-way $K$-shot question.
3.2 Few-Shot Detector
Our FSD is based on the prototypical network (Snell et al., 2017), an effective approach for addressing the few-shot question above. FSD aims at learning a metric space in which samples from the same class are close in distance, while samples from different classes are relatively distant from each other. Thus, we can represent each class by one single vector, known as the prototypical representation. For each sample in the query set, classification is then conducted by simply finding the nearest representation to it.
In detail, as depicted in Fig. 1, FSD first adopts a neural network $f_{\phi}:\mathbb{R}^{D}\to\mathbb{R}^{M}$ to compute the $M$-dimensional vector in the metric space of an image, where $\phi$ is the learnable parameters. A prototypical representation $\mathbf{c}_{i}$ is defined as the mean of embedded vectors from its support set $\mathcal{S}_{i}$:
$\mathbf{c}_{i}=\frac{1}{|\mathcal{S}_{i}|}\sum_{\mathbf{x}_{j}\in\mathcal{S}_{i}}f_{\phi}(\mathbf{x}_{j})$ (1)
After calculating all the $N$ prototypical representations, for an image $\mathbf{x}$ in the query set, we compute its vector $f_{\phi}(\mathbf{x})$ and the distances between it and all the representations $d(f_{\phi}(\mathbf{x}),\mathbf{c}_{i})$, where $d:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}_{0}^{+}$ is a distance function. The probability that the query sample $\mathbf{x}$ belongs to class $i$ is calculated by:
$p(y=i|\mathbf{x})=\mathrm{softmax}_{1\leq i\leq N}(-d(f_{\phi}(\mathbf{x}),\mathbf{c}_{i}))$ (2)
The image $\mathbf{x}$ is then assigned to the class with maximum probability, or the nearest prototypical representation to it.
During training, we randomly selected $N_{c}$ classes from the training set $\mathcal{D}_{train}$. For each selected class $i$, we randomly choose $N_{s}$ samples as support set $\mathcal{S}_{i}$ and another $N_{q}$ samples as query set $\mathcal{Q}_{i}$. We compute the $N_{c}$ prototype centers $\mathbf{c}_{i}$ according to equation 1. The prototypical network outputs a probability distribution over the $N_{c}$ classes for each query sample $\mathbf{x}$ in query set. The optimization target is to minimize the the negative log-probability $J(\phi)$. Pseudo code to compute the loss $J(\phi)$ for a training step is provided in Algorithm 1.
Algorithm 1 Prototypical Loss Computation Algorithm
  Input: Training set $\mathcal{D}=\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\cdots\mathcal{D}_{N}$, where $\mathcal{D}_{k}$ denotes the subset of class $k$, number of training class $N_{C}$, number of support examples per class $N_{S}$, number of query examples per class $N_{Q}$.
  Output: training loss $J$ of one step.
  $V\leftarrow\textit{randomly select }N_{c}\textit{ classes from }\{1,2,\ldots,N\}$
  for $k$ in $V$ do
     $S_{k}\leftarrow\textit{randomly select }N_{s}\textit{ samples from }\mathcal{D}_{k}$
     $Q_{k}\leftarrow\textit{randomly select }N_{q}\textit{ samples from }\mathcal{D}_{k}/S_{k}$
     $\mathbf{c}_{k}\leftarrow\frac{1}{N_{s}}\sum_{\mathbf{x}_{i}\in S_{k}}f_{\phi}(\mathbf{x}_{i})$
  end for
  $J\leftarrow 0$
  for $k$ in $V$ do
     $J\leftarrow J+\frac{1}{N_{c}N_{q}}\sum_{\mathbf{x}_{j}\in Q_{k}}\log\mathrm{softmax}(-d(f_{\phi}(\mathbf{x}_{j}),\mathbf{c}_{k}))$
  end for
  Return $J$
3.3 Zero-shot Classification
We propose a zero-shot detection way to verify the applicability of FSD in real-world settings. In the zero-shot task, there is no additional samples from the test classes. Instead, we are provided with a meta-data vector for each class in the training set to categorize the unseen images. We place considerable emphasis on zero-shot classification, since in practical scenarios, it is unfeasible to obtain test samples from every possible generative model. The zero-shot results can represent the performance of our model in general scenarios.
In this work, we simply define the meta-data vector as the prototypical representation derived from an extensive collection of samples. That is, we randomly select plenty of images from each class in the training set and then compute their prototypical representations following Equation 1. Once the nearest representation of a test image is labeled as a synthetic class, it will be regarded as a fake image.
4 Experiment
Table 1: Comparison of our FSD and previous studies. We train several classifiers on different subsets and report the accuracy for each unseen test subset. The best results are highlighted in boldface, while the second-optimal results are marked with underline.
Method Test Subset Avg
Midjourney GLIDE ADM SD VQDM BigGAN Acc(%)
CNNSpot (Wang et al., 2020) 58.2 57.1 57.0 70.3 56.7 56.6 59.3
Spec (Zhang et al., 2019b) 56.7 65.4 57.9 72.4 61.7 64.3 63.1
F3Net (Qian et al., 2020) 55.1 57.8 66.5 73.1 62.1 56.5 61.9
GramNet (Liu et al., 2020) 58.1 65.3 58.7 72.8 57.8 61.2 62.3
DIRE (Wang et al., 2023) 65.0 69.1 61.9 73.7 63.4 56.7 65.0
LARE2 (Luo et al., 2024) 66.4 81.3 66.7 87.3 84.4 74.0 76.7
FSD (zero-shot) 75.1 93.9 74.1 88.0 69.1 62.1 77.1
FSD (10-shot) 80.9 97.1 79.2 88.8 76.2 82.2 84.1
In this section, we first introduce the experimental setups and then provide a comparison between FSD and previous studies. We also visualize the feature space to highlight the superiority of our approach.
4.1 Datasets and Evaluation Metrics
Datasets. We evaluate our proposed method on the widely used GenImage dataset (Zhu et al., 2023), which contains $1,331,167$ real images from ImageNet (Deng et al., 2009) and $1,350,000$ generated images from $7$ diffusion models and one GAN. The diffusion generators include Midjourney (mid, 2022), Stable Diffusion V1.4 (Rombach et al., 2022), Stable Diffusion V1.5 (Rombach et al., 2022), Wukong (wuk, 2022), ADM (Dhariwal & Nichol, 2021), GLIDE (Nichol et al., 2022) and VQDM (Gu et al., 2022). And the only GAN is BigGAN (Brock et al., 2019). The real images are first divided into $8$ subsets. And each subset is then further partitioned into training part and test part. Each generator is paired with a specific subset and uses the category labels of the real images within the subset to produce synthetic images. Therefore, GenImage dataset consists of $8$ fake-vs-real subsets corresponding to the $8$ generative models.
In this study, the synthetic images in different subsets are categorized into distinct classes, whereas the real images across all subsets are classified into a single class. We also notice that $3$ diffusion models SD v1.4, SD v1.5 and Wukong share an identical model structure and it is difficult to distinguish among them. Thus, we combine their corresponding subsets into a unified subset, named Stable Diffusion (SD). Finally, the dataset used in our experiments comprises $7$ classes of images from different sources.
Evaluation metrics. Following previous AI-generated image detection studies (Zhu et al., 2023; Wang et al., 2023; Luo et al., 2024), we adopt the accuracy (ACC) and average precision (AP) as our evaluation metrics. The threshold step for computing AP is set to $0.1$.
4.2 Implementation Details
To be more comparable with previous works, we adopt the ResNet-$50$ (He et al., 2016) pretrained on ImageNet (Deng et al., 2009) as the backbone of our model, which outputs a prototype vector of 1024 dimensions. Following Wang et al. (2023), the input images are first resized to $256\times 256$ and then randomly cropped to $224\times 224$ with a random horizontal flip during training. In contrast, only a center crop to $224\times 224$ is performed after resizing the images during testing. The distance in the metric space is measured with squared Euclidean distance.
To fit the classification task, we apply different few-shot manners to training and testing. In detail, we select one synthetic subset as the test data and utilize the rest $6$ subsets as the training data for one classifier. At every training step, $3$ classes are randomly selected from the training set, with $5$ random samples chosen from each class for the support batch and another $5$ samples for the query batch. We employ Adam as the optimizer to minimize the cross entropy loss with a base learning rate of $10^{-4}$. We also adopt a StepLR scheduler with $\gamma=0.5$ and $step=80000$. The training of each classifier is carried out for $200,000$ steps with a batch size of $16$. During testing, we conduct a $2$-way classification task by comparing the selected test class against the real class. For the zero-shot detection, we randomly select $1024$ samples from each class in the training set to compute the meta-data vectors. Our method is implemented by PyTorch library and all the experiments are conducted on a single A100 with 40GB memory.
Table 2: Result of $10$-shot cross-generator image classification on different subsets. Each classifier is trained on the dataset excluding the subset from the first column. All classifiers are tested on the $6$ synthetic classes, reporting Accuracy / Average Precision (%).
Excluding Subset Test Subset
Midjourney GLIDE ADM SD VQDM BigGAN
Midjourney 80.9 / 84.6 99.9 / 99.9 98.5 / 99.3 97.1 / 98.7 99.5 / 99.9 88.0 / 92.9
GLIDE 86.8 / 89.9 97.1 / 98.0 97.9 / 98.9 97.1 / 98.8 99.2 / 99.7 91.9 / 97.1
ADM 87.6 / 91.8 99.8 / 99.9 79.2 / 83.8 94.8 / 97.2 98.8 / 99.4 91.0 / 96.1
SD 86.1 / 89.7 99.9 / 99.9 97.4 / 98.8 88.8 / 92.5 96.6 / 98.5 89.5 / 95.4
VQDM 82.4 / 85.9 99.9 / 99.9 97.3 / 98.6 95.6 / 98.0 76.2 / 79.4 83.5 / 89.1
BigGAN 88.9 / 91.6 99.9 / 99.9 98.3 / 99.3 98.1 / 99.3 96.4 / 98.3 82.2 / 86.8
4.3 Comparison to Existing Methods
We compare FSD with several state-of-the-art synthetic image detection methods, including CNNSpot (Wang et al., 2020), Spec (Zhang et al., 2019b), F3Net (Qian et al., 2020), GramNet (Liu et al., 2020), DIRE (Wang et al., 2023) and LARE2 (Luo et al., 2024). The first four methods are model-agnostic, focusing on identifying synthetic artifacts within fake images. And the last two methods leverage the reconstruction error of Stable Diffusion to differentiate between fake and real images. To conduct a comprehensive comparison, we train $6$ classifiers on different training sets, each of which excludes a specific test subset. However, previous studies typically provide several classifiers, and each classifier is trained on one distinct subset and evaluated on the others. To minimize the difference between our method and previous researches, we report the average performance of those classifiers over each test subset in previous studies, which can better represent their generalization ability on each class. The results are summarized in Table 1.
As illustrated in Table 1, our FSD attains state-of-the-art performance across $5$ out of the $6$ test classes, surpassing previous state-of-the-art by $+7.4\%$ accuracy in average. The suboptimal performance of our model on the VQDM class can be attributed to the involvement of image quantization in VQDM, which significantly differs from the other generative models in our experiments. Another reason is that LARE2 is specifically designed to detect synthetic images generated by diffusion models, which enables it to achieve unexpectedly high performance on the VQDM class. Although our few-shot method may not be directly comparable to other approaches due to the accessing of the few samples in the test class, we still regard this discrepancy as a significant strength of our model. The definite improvement in performance demonstrates FSD’s capability to effectively leverage a limited number of unseen samples.
Based on Table 1, we find that FSD achieves notable performance in zero-shot scenarios, in which the prototypical representations are calculated from samples in the training classes. The result shows that our approach is also suitable for detecting synthetic images even when samples from the same sources are not provided, thus broadening the application scope of FSD. We owe this advantage to FSD’s capability to capture not only the intra-class characteristics but also the common features across all classes. Another observation is that the discrepancy between zero-shot results and few-shot results remains notable, which indicates the necessity of gathering training data from a wider range of generative models.
4.4 Cross-Generator Classification
We evaluate the $10$-shot performance of FSD on both seen classes and unseen classes. We also train $6$ classifiers for the cross-generator image classification task, in which each classifier is trained on $6$ out of $7$ subsets in the dataset and subsequently tested on the remaining one. We report the accuracy and average precision across each scenario and summarize the results in Table 2.
As shown in Table 2, each row presents the classification results for one classifier across all the $6$ synthetic classes. The results show that it is easy for a classifier to detect fake images from those classes encountered during training, which is consistent with the previous studies (Zhu et al., 2023; Luo et al., 2024). When it comes to detecting the unseen classes, as shown in the diagonal of Table 2, FSD suffers from a slight decline in performance, demonstrating the generalization capability of our model. Another finding is that FSD performs nearly perfectly on the VQDM class when this class has been included in the training set. However, the performance significantly declines when this class has been excluded. This fact shows that it is still challenging to generalize across unseen generative models, indicating that our approach remains inadequate in capturing the discriminative features between classes.
We also observe the varying performance of FSD across different classes. For instance, images generated by GLIDE, ADM, SD and VQDM are relatively easy to detect by our model when these classes are included in the training set, while those from Midjourney and BigGAN are a little challenging to identify. Particularly, images from the GLIDE class can be distinguished with exceptional accuracy, even when not included in the training set. These facts indicate the necessity of treating images from different sources as distinct categories, which has not been fully explored in previous researches. Moreover, it is essential to collect images from representative generative models for training, as generalization on similar generators is easier than on those that are vastly different.
(a) Test on Midjourney
(b) Test on GLIDE
(c) Test on ADM
(d) Test on Stable Diffusion
(e) Test on VQDM
(f) Test on BigGAN
Figure 3: Visualization of network features for different classifiers over different classes using t-SNE. Each classifier is trained on those subsets excluding the test one in each sub-figure. The features are extracted from the final layer of our model.
4.5 Visualization
To further analyze the effectiveness of FSD, we utilize t-SNE (van der Maaten & Hinton, 2008) visualization to illustrate the feature space of classifiers on different training sets, as shown in Figure 3. Each classifier is trained on those subsets excluding the test one in each sub-figure. We randomly selected $1024$ samples from the test part of each class for visualization. And the features are extracted from the final layer of our model.
From Figure 3, we can see that features of images from different generative models exhibit distinct distributions, each of which is marked by a unique color. Most classes within the training set can be well-separated due to the large margins among their distributions. However, the boundary of the Midjourney class represented in red is not as clear as that of the others. This can be attributed to its API service, which may utilize multiple generative models for image generation based on our inference. Although the boundary of the test class in each sub-figure is a little ambiguous, the distribution of this class still remains identifiable, which demonstrates that our method successfully captures the intra-category features of unseen data.
The results also support the validity of our approach that categorizes samples from different generators as different classes. However, such categorization also introduces the issue of eliminating the significant differences between seen data and unseen data. As illustrated in Figure 3(e), The distribution of the VQDM class in orange exhibits a large overlap with that of the real class in gray, due to the great difference in image tokenization between VQDM and other generative models. This also indicates that our method still faces the challenge in detecting images from models that are significantly different from those generative models used in training, which will be a main focus of our future work.
4.6 Ablation Study
Impact of the number of shots. An important aspect is to determine the minimum number of samples required by our approach to achieve satisfactory performance. We choose different numbers of support samples in each class, ranging from $1$ to $200$. To present a comprehensive overview of this aspect, we conducted the experiment across all the $6$ classifers described in Section 4.4. We evaluate each classifier on the corresponding test class with varying numbers of support samples. To maintain a consistent total number of test samples across different settings, We set a fixed ratio of $1:3$ between the support set and the query set. The results are shown in Figure 4, in which each fold line represents the performance of a classifier across varying sizes of support set.
From Figure 4, we can see that using more support samples results in better performance. However, it also increases the required computational resources. We find that detecting with $10$ shots achieves an optimal balance between performance and resource consumption for most of the test classes. For example, our model brings the accuracy on the ADM class from $62.6\%$ to $79.2\%$ when increasing the number of shots from $1$ to $10$, with an improvement of $+16.6\%$. However, it achieves $81.7\%$ accuracy when using $200$ shots, gaining only an increase of $+2.5\%$ over the $10$-shots scenario. As it is easy to collect $10$ samples from online generative models and the computation consumption is very little, We ultimately decide to use the $10$-shots results of FSD for comparison.
1111333355551010101025252525505050501001001001002002002002005050505060606060707070708080808090909090100100100100Number of shots per classAccuracy (%)MidjourneyGLIDEADMSDVQDMBigGAN
Figure 4: Influences of the number of shots. Each classifier is trained on those subsets excluding the test one. And $10$-shots is the most cost-effective setting across all the categories.
(a) FSD
(b) Binary Classifier
Figure 5: Visualization of network features for our FSD (a) and a binary classifier (b) on ADM class. In the feature space of FSD, images from the unseen class in purple are tightly clustered together. However, they are much more dispersed in the feature space of the binary classifier.
Impact of the multi-class classification. Different from previous studies which treat the fake images as a single class and train an binary classifier to detect them, we category the images from different generators as distinct classes. We argue that distinguishing images from different generators can also benefit the detection performance. Thus, we train another binary classifier and compare it with our method on the ADM class, which is difficult to detect according to our experiments. Both classifiers utilize the ResNet-$50$ as backbone and are trained on the dataset excluding the ADM class. We also apply the t-SNE visualization and random select $1024$ samples for visualizing. Features of the binary classifier are extracted from the output layer of the ResNet-$50$, with a dimension of $2048$. The visualization results are shown in Figure 5.
Figure 5(a) shows that our FSD successfully distinguishes those images across different classes which has been used in training. Images from the unseen ADM class dyed in purple also cluster together in the visualization picture. In contrast, the traditional binary classifier fails to distinguish images from Midjourney and Stable Diffsion, as shown in Figure 5(b). And the unseen samples in purple are quite dispersed, failing to form a compact cluster. Thus, distinguishing them from the real ones can be extremely difficult. This fact indicates that our method is able to extract intra-category common features from unseen classes, thereby enabling an accurate differentiation among various classes. Moreover, we find that it is difficult for a classifier to distinguish samples from Stable Diffusion V1.4 and Stable Diffusion V1.5. One plausible explanation could be the shared usage of an identical backbone and VAE, which makes their products significantly similar. This implies that we should carefully select the representative generative models to build the training set, which We leave for future work.
5 Conclusion and Limitations
The key point of detecting fake images in real-world scenarios is to make full use of the limited images from unseen generative models. In this paper, we propose Few-Shot Detector (FSD), a novel AI-generated image detector based on few-shot learning. Experiments show that FSD achieves state-of-the-art performance in synthetic image detection with only $10$ samples from the test generator, demonstrating the strong generalization ability of our model. We visualize the feature space to highlight the advantage of distinguishing images from different sources, which has not been widely adopted among previous studies.
However, our method still suffers from a few limitations. Since it is not feasible to obtain few-shot samples from all generative models, our collection of synthetic images is confined to part of them, which may lead to a decrease in detection performance. Another weakness is that our method does not perform satisfactorily when detecting images from a vastly different generative model. We aim to address these issues in our future work.
Acknowledgments
This work is supported by the National Science and Technology Major Project (No. 2022ZD0116314) and the National Science Foundation of China (No. 62106249).
Impact Statement
AI-generated image detection has been a subject of extensive researches and is regarded as an important technology to reduce cheating, misinformation and harm on internet platforms and social media. Our work aims at enhancing fake image detection performance and achieving better generalization with a few samples, rather than increasing synthetic data. All the images used in our experiment are from open-source dataset. We believe our work can contribute to the development of deepfake detecting applications in many real-world scenarios.

### v2

@@
 Current fake image detectors trained on large synthetic image datasets perform satisfactorily on limited studied generative models.
-However, they suffer a notable performance decline over unseen models.
+However, these detectors suffer a notable performance decline over unseen models.
 Besides, collecting adequate training data from online generative models is often expensive or infeasible.
-To overcome these issues, we propose Few-Shot Detector (FSD), a novel AI-generated image detector which learns a specialized metric space to effectively distinguish unseen fake images by utilizing very few samples.
-Experiments show FSD achieves state-of-the-art performance by $+7.4\%$ average ACC on GenImage dataset.
-More importantly, our method is better capable of capturing the intra-category common features in unseen images without further training.
+To overcome these issues, we propose Few-Shot Detector (FSD), a novel AI-generated image detector which learns a specialized metric space for effectively distinguishing unseen fake images using very few samples.
+Experiments show that FSD achieves state-of-the-art performance by $+11.6\%$ average accuracy on the GenImage dataset with only $10$ additional samples.
+More importantly, our method is better capable of capturing the intra-category commonality in unseen images without further training.
+Our code is available at https://github.com/teheperinko541/Few-Shot-AIGI-Detector.
 Deepfake Detection, Synthetic Image Detection, Few-Shot Learning
 1 Introduction
-The development of generative models has led to significant strides in synthesizing photorealistic images, making it much difficult to distinguish AI-generated images from real ones.
-Over the past few years, diffusion-based models (Ho et al., 2020) have exceeded GANs (Goodfellow et al., 2014) and become the main stream paradigm of image generation, due to their strong ability of synthesizing images of high-quality.
-With the rapid release of open-source models and web APIs, users can easily create unrealistic images with brief texts.
-However, the increasing abuse of deepfake technology has dramatically influenced the human society and sparked unprecedented concerns on the credibility of online information.
-To curb the spread of AI-generated images used for malicious purposes, there is a growing urgency to develop detection methods for distinguishing fake images from authentic ones.
-Early works (Zhang et al., 2019b; Chai et al., 2020) focus on identifying artifacts in images, which prove to be particularly useful for detecting GAN-generated images.
-With the improvement of synthetic image datasets (Zhu et al., 2023; Boychev & Cholakov, 2024), detecting fake images from seen models is no longer difficult.
-Thus, a new challenge is to develop a classifier that generalizes across diverse unseen generative models.
+The development of generative models has led to significant strides in synthesizing photorealistic images, making it much more difficult to distinguish AI-generated images from real ones.
+Over the past few years, diffusion-based models (Ho et al., 2020) have exceeded GANs (Goodfellow et al., 2014) and become the mainstream paradigm of image generation, due to their exceptional ability to synthesize high-quality images.
+With the rapid release of open-source models and web APIs, users can easily create vivid images from brief textual descriptions.
+However, the increasing abuse of deepfake technology has dramatically influenced human society and sparked unprecedented concerns on the credibility of online information.
+To curb the spread of AI-generated images used for malicious purposes, there is a growing urgency to develop detection methods capable of distinguishing fake images from authentic ones.
+Early works (Zhang et al., 2019b; Chai et al., 2020) focus on identifying artifacts in synthetic images, which have proved to be particularly useful for detecting GAN-generated images.
+With the improvement of synthetic image datasets (Zhu et al., 2023; Boychev & Cholakov, 2024), detecting fake images from seen generative models is no longer a difficult task.
+Thus, a novel challenge arises in developing classifiers that generalize across unseen models.
 To address this problem, some studies (Ojha et al., 2023; Liu et al., 2024) adopt a large vision-language model to analyze images from a semantic perspective, while others (Wang et al., 2023; Luo et al., 2024) leverage the reconstruction ability of diffusion models.
 Another difficulty is the acquisition of training data.
-It is often expensive or infeasible to collect enough images from closed-source models such as DALL-E (Ramesh et al., 2021) and Midjourney (mid, 2022), due to their service prices and access restrictions.
-However, most existing fake image detectors (Wang et al., 2023; Ojha et al., 2023) require large amounts of training data to achieve good generalization, while fine-tuning without adequate data often results in overfitting.
-Furthermore, the ongoing release and updating of generative models pose a significant challenge for any trained detectors to keep up.
-After gaining a deep insight into these limitations, we propose a distinctive scheme based on few-shot learning to overcome them and design a comprehensive model to validate the efficacy of our solution.
+It is often expensive or infeasible to collect enough images from closed-source models such as DALL-E (Ramesh et al., 2021) and Midjourney (mid, 2022).
+This is due to their service prices and access restrictions.
+However, most existing fake image detectors (Wang et al., 2023; Ojha et al., 2023) require large amounts of training data to achieve good generalization, and fine-tuning without adequate data often results in overfitting.
+Furthermore, the ongoing release and updating of generative models pose a significant challenge for existing detection systems to keep up.
+After gaining a deep insight into these limitations, we propose an innovative approach based on few-shot learning to overcome them.
+A robust model is also developed to validate the efficacy of our solution.
 We first point out that current AI-generated image detection is a domain generalization task.
-Previous studies have been dedicated to finding a common indicator effective for all generative images.
+Previous studies have been dedicated to discovering a universal indicator effective for detecting diverse fake images.
 ̵However, they overlook the significant distinctions among data from different domains.
 We observe that images from unseen domains can actually be obtained in many real-world scenarios.
-Based on this fact, the complicated task can be transformed to a slightly easier one, called few-shot classification, by using relative few samples from unseen domains.
-Thus, we can extract rich domain information from these samples and use it to generalize across the unseen data.
-Based on the above analysis, we propose Few-Shot Detector (FSD), a novel method which can detect fake images with few given samples.
-In the few-shot classification task, synthetic images from different generative models are classified into separate categories, while the real images are classified into one class.
-The goal is to train a classifier capable of generalizing to unseen classes not in the training set, with only a small number of examples from them.
-To solve this problem, FSD utilizes a prototypical network (Snell et al., 2017), which learns a metric space to be able to handle unseen data.
-Different from end-to-end classification methods, FSD requires receiving a few samples from several classes to compute the corresponding prototypical representations.
-The test image is then classified with the nearest-neighbor method.
-Thus, the bias in unseen classes which can not be learned from the training set can now be automatically applied to the classifier without extra training.
-Experiments show that FSD obtains notable improvements over the state-of-the-art method by $+7.4\%$ average accuracy on GenImage dataset.
-We observe that only $10$ samples from the unseen classes can significantly improve the detection performance, which shows great privilege of our method.
-We also show that our approach is better equipped to keep pace with the rapid advancements in image generation, thereby highlighting the effectiveness of categorizing generated images by their sources.
-Our main contributions are as follows:
-  • To the best of our knowledge, we are the first to reconceptualize AI-generated image detection as a few-shot classification task, bringing it closer to the real-world scenario applications.
-  • We introduce an innovative synthetic image detection method, named Few-Shot Detector (FSD). Our approach is able to utilize few samples from unseen domains to achieve better detection performance.
+Based on this fact, the complicated task can be transformed into a more tractable one, called few-shot classification, through limited samples from unseen domains.
+Thus, we can extract rich domain-specific features from the given samples and use them to generalize across the unseen data, as shown in Figure 1.
+Figure 1: Challenge facing previous works and our solution.
+Most classifiers exhibit significant performance degradation over unseen data.
+We address this challenge by introducing a few-shot strategy which is able to make full use of the limited unseen samples.
+Based on the above analysis, we propose Few-Shot Detector (FSD), a novel method which is able to detect fake images with few given samples.
+In the few-shot classification task, synthetic images from different generative models are classified into separate categories, whereas the real images are classified into a single category.
+The objective is to develop a classifier capable of generalizing to unseen classes that are not in the training set, with only a limited number of examples from them.
+To solve this problem, FSD utilizes the Prototypical Network (Snell et al., 2017), which learns a metric space to effectively handle unseen data.
+The given samples are used to compute the prototypical representations of different classes.
+The test image is then classified using the nearest-neighbor method by comparing it against these representations.
+Thus, FSD can automatically adjust for bias in previously unseen classes without additional training or fine-tuning.
+Experiments show that FSD achieves notable improvements over the state-of-the-art method by $+11.6\%$ average accuracy on the GenImage dataset.
+We observe that only $10$ samples from the tested classes can significantly improve the detection performance, which highlights the superiority and practicality of FSD.
+We also show that our approach is better equipped to keep pace with the rapid advancements in image generation, thereby demonstrating the effectiveness of categorizing generated images by their sources.
+Our contributions are summarized as follows:
+  • To the best of our knowledge, we are the first to reconceptualize AI-generated image detection as a few-shot classification task, bringing this task closer to real-world applications.
+  • We introduce an innovative synthetic image detection method, named Few-Shot Detector (FSD).
+Our approach is able to utilize a few samples from unseen domains to achieve better detection performance.
   • Our experiments demonstrate the generalization ability of FSD, which significantly outperforms current state-of-the-art methods and can easily deal with new generative models without further training.
@@
 Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) have revolutionized the field of generative modeling over the last decade.
-By simultaneously training a generator and a discriminator in an adversarial manner, GANs can generate samples of high resolution and good quality (Karras et al., 2020; Brock et al., 2019).
-However, they are prone to mode collapse, making it difficult for them to capture the full data distribution.
-An alternative technical route is likelihood-based modeling, which estimates the parameters of data distribution by maximizing the likelihood function.
-Within this framework, variational autoencoders (VAEs) (Kingma & Welling, 2014) ensure a smooth latent space which facilitates interpolation and exploration of new samples.
-Meanwhile, normalizing flow models (Papamakarios et al., 2021) allow for exact computation of the log-likelihood, which enables them to capture more complex data distributions.
-However, neither of them can match the high-quality image generation capabilities of GANs.
-Diffusion Models (Ho et al., 2020; Song et al., 2021) have recently achieved remarkable performances in image synthesis.
-By reversing a gradual noising process, diffusion models can generate an image from random noise with several denoising steps.
-Ho & Salimans (2022) introduce a classifier-free guidance that enables the generation of images from text descriptions, eliminating the need for training a separate classifier.
-Another significant advancement is latent diffusion model (LDM) (Rombach et al., 2022), which applies diffusion process in latent space with a powerful autoencoder.
+By simultaneously training a generator and a discriminator in an adversarial manner, GANs can generate samples of high resolution and good quality (Brock et al., 2019; Karras et al., 2020).
+However, they are prone to mode collapse, which makes it difficult for them to fully capture the data distribution.
+An alternative approach is likelihood-based modeling, which estimates the parameters of data distributions by maximizing the likelihood function.
+Based on this methodology, variational autoencoders (VAEs) (Kingma & Welling, 2014) ensure a smooth latent space that facilitates interpolation and exploration of new samples.
+In contrast, normalizing flow models (Papamakarios et al., 2021) enable exact log-likelihood computation, allowing them to capture more complex data distributions.
+However, neither of them can match GANs in high-quality image generation.
+Diffusion Models (Ho et al., 2020; Song et al., 2021) have recently achieved remarkable performance in image synthesis.
+By reversing a gradual noising process, diffusion models can generate images from random noise through successive denoising steps.
+Ho & Salimans (2022) introduce classifier-free guidance which enables generating images from text descriptions, eliminating the need for training a separate classifier.
+Another significant advancement is the Latent Diffusion Model (LDM) (Rombach et al., 2022), which applies the diffusion process in latent space with a powerful autoencoder.
 This approach greatly enhances the visual fidelity of synthetic images while substantially reducing the required computational resources.
-Based on vast amounts of training data and substantial computing resources, large latent diffusion models (Podell et al., 2024; Esser et al., 2024) have become the most potent technique in image generation and the main source of synthetic images at present.
-Figure 1: Detection pipeline of Few Shot Detector.
-FSD first maps the real and fake images from support set to the metric space by prototypical network and calculates the representation of each type.
-A test image is then mapped into the same space and classified according to the principle of nearest neighbor.
+Leveraging vast amounts of training data, large-scale latent diffusion models (Podell et al., 2024; Esser et al., 2024) have emerged as the leading technique in image generation and the main source of synthetic images at present.
+Figure 2: Detection pipeline of Few-Shot Detector.
+FSD first maps the real and fake images from the support set to the metric space by the Prototypical Network and calculates the representation of each type.
+A test image is then mapped into the same space and classified according to the nearest-neighbor principle.
 2.2 AI-generated Image Detection
 Model-agnostic detection method.
-Learning based methods have been applied to detecting fake images from GANs (Zhang et al., 2019b; Qian et al., 2020; Chai et al., 2020).
-They treat synthetic image detection as a binary classification task, aiming to find the synthetic clues in generated images.
-Many studies (Zhang et al., 2019a; Gragnaniello et al., 2021) have found that classifiers do not generalize over unseen generative models, bringing generalization ability to the primary research objective.
-Meanwhile, Wang et al.
+Learning-based methods have been applied to detecting fake images from GANs (Zhang et al., 2019b; Qian et al., 2020; Chai et al., 2020).
+They treat synthetic image detection as a binary classification task, aiming to find synthetic clues in generated images.
+Numerous studies (Zhang et al., 2019a; Gragnaniello et al., 2021) have found that classifiers often fail to generalize across unseen generative models, bringing generalization ability to the primary research objective.
+Wang et al.
 (2020) reveal that detectors for CNN-generated images can be surprisingly generalizable due to the common systematic flaws shared among CNNs.
@@
 (2023) analyze multi-view features to learn a robust representation, while Tan et al.
-(2024b) focus on high-frequency information to learn the source-agnostic features.
+(2024b) focus on high-frequency information to capture source-agnostic features.
 However, these common cues disappear when it comes to diffusion models, leading to the poor performance of earlier methods when applied to them.
@@
 For instance, Ojha et al.
-(2023) prove that the informative image features from a pretrained CLIP model (Radford et al., 2021) can successfully be used for detecting fake images.
+(2023) prove that the informative image features from a pretrained CLIP model (Radford et al., 2021) can be successfully used for detecting fake images.
 Other studies (Liu et al., 2024; Tan et al., 2024a) deeply integrate textual information with images, thereby further enhancing the generalization capability.
-Moreover, patch-based detection method (Chen et al., 2024) has been proposed to reduce computational resources and accelerate detection speed.
-This method extracts features from a few patches instead of the whole image, enabling more efficient discrimination between fake and real images.
+Moreover, a patch-based detection method (Chen et al., 2024) extracts features from a single patch rather than the entire image, which significantly accelerates the detection speed.
 Diffusion-based detection method.
-The capability of open-source diffusion models to reconstruct images has opened a new technological route.
-DIRE (Wang et al., 2023) first uses the reconstruction loss to train a classifier, based on the observation that synthetic images can be reconstructed better than real ones.
-Following this discovery, SeDID (Ma et al., 2023) analyzes the loss errors between noised and denoised features at specific steps of processing, while LARE2 (Luo et al., 2024) focuses on the single-step reconstruction result in latent space.
-Both of them are able to improve the detection speed by not fully completing the entire reconstruction process.
-FakeInversion (Cazenavette et al., 2024) uses a BLIP (Li et al., 2022) to obtain a text description of the image for better reconstruction, which is capable of detecting images from unseen high-fidelity generators.
-AEROBLADE (Ricker et al., 2024) evaluates the autoencoder (AE) in different latent diffusion models and presents a training-free method for detecting fake images by AE reconstruction errors.
-These approaches achieve great success on open-source generative models.
-However, the differences between various diffusion models still can not be ignored, which makes it particularly challenging for them to detect synthetic images produced by numerous generative models, especially those that are closed-source.
+The reconstruction capability of open-source diffusion models has opened a new technological pathway.
+DIRE (Wang et al., 2023) first uses the reconstruction loss to train a classifier, based on the observation that synthetic images can be reconstructed with higher fidelity compared to real ones.
+SeDID (Ma et al., 2023) analyzes the loss errors between noised and denoised features at specific steps of processing, while LARE2 (Luo et al., 2024) focuses on the single-step reconstruction result in latent space.
+Both of them improve the detection speed by not fully completing the entire reconstruction process.
+FakeInversion (Cazenavette et al., 2024) employs BLIP (Li et al., 2022) to obtain a text description of the image for better reconstruction, which is capable of detecting unseen high-fidelity generated images.
+AEROBLADE (Ricker et al., 2024) evaluates autoencoder (AE) in different latent diffusion models and presents a training-free method for detecting fake images based on AE reconstruction errors.
+Although these approaches have achieved great success on open-source generative models, they still face challenges in detecting synthetic images from various inaccessible models due to the substantial differences among them.
 3 Method
+In this section, we first describe the few-shot classification task for synthetic image detection.
+Then we elaborate on our novel representation, namely Few-Shot Detector (FSD), as illustrated in Figure 2.
+3.1 Few-shot Synthetic Image Detection
+Few-shot classification is a specialized machine learning task, where the classifier is challenged to generalize to entirely new classes not present in the training set, using only a limited number of examples from these novel classes.
+Different from end-to-end detection methods that aggregate all synthetic images produced by different generative models into a single fake class, we classify these images based on their generators.
+This means that we consider the productions of each model as an individual class.
+Conversely, authentic images continue to be grouped into a single class.
+During detection, the support set comprises labeled images aiding in the classification task, while the query set contains unlabeled images to be classified.
+The objective of this task is to determine the category of each image in the query set based on the support set.
+In detail, we are given a support set $\mathcal{S}=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\cdots\cup\mathcal{S}_{N}$, which comprises samples from $N$ new classes.
+Each subset $\mathcal{S}_{i}=\{(\mathbf{x}_{i1},y_{i1}),(\mathbf{x}_{i2},y_{i2}),\ldots,(\mathbf{x}_{iK},y_{iK})\}$ denotes the collection of images from class $i$.
+$K$ is the number of samples in each class.
+$\mathbf{x}_{ij}\in\mathbb{R}^{D}$ is an image of $D$ dimensions and $y_{ij}=i$ indicates its corresponding label.
+Each unlabeled image in the query set will be assigned to one of the $N$ classes.
+This is the so-called $N$-way $K$-shot task.
 (a) training
 (b) testing
-Figure 2: Prototypical network process.
-During training, the optimal center of each class is learned by minimizing the distances between query samples and their corresponding centers.
+Figure 3: Prototypical calculation.
+During training, the metric space is learned by minimizing the distances between query samples and their corresponding centers.
 During testing, the test image is only classified between real and fake.
-In this section, we first describe the few-shot classification task of synthetic image detection.
-Then we elaborate on our novel representation, namely Few-Shot Detector (FSD), as illustrated in Fig.
-1.
-3.1 Few-Shot Synthetic Image Detection
-Few-shot classification is a specialized machine learning task, in which the classifier is challenged to generalize to entirely new classes that are not present in the training set, using only a limited number of examples from these novel classes.
-Different from previous studies that categorize all the synthetic images produced by different generative models into a single fake class, we classify the images based on their generators.
-That is, we treat each group of them as a distinct class.
-In contrast, all the authentic images are classified into a single class.
-During detection, the set of additional labeled images used to support the classification task is known as the support set, while the set of those to be tested is referred to as the query set.
-The goal of this task is to determine which category each image in the query set belongs to, based on the support set.
-In detail, we are given a support set comprising samples from $N$ new classes $\mathcal{S}=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\cdots\cup\mathcal{S}_{N}$, where $\mathcal{S}_{i}=\{(\mathbf{x}_{i1},y_{i1}),(\mathbf{x}_{i2},y_{i2}),\ldots,(\mathbf{x}_{iK},y_{iK})\}$ is the image subset of class $i$.
-$K$ is the number of samples in each class.
-$\mathbf{x}_{ij}\in\mathbb{R}^{D}$ is an image feature of $D$ dimension and $y_{ij}=i$ is the corresponding label.
-Each unlabeled image in the query set will be assigned to one of the $N$ classes.
-And this is the so-called $N$-way $K$-shot question.
 3.2 Few-Shot Detector
-Our FSD is based on the prototypical network (Snell et al., 2017), an effective approach for addressing the few-shot question above.
-FSD aims at learning a metric space in which samples from the same class are close in distance, while samples from different classes are relatively distant from each other.
-Thus, we can represent each class by one single vector, known as the prototypical representation.
-For each sample in the query set, classification is then conducted by simply finding the nearest representation to it.
-In detail, as depicted in Fig.
-1, FSD first adopts a neural network $f_{\phi}:\mathbb{R}^{D}\to\mathbb{R}^{M}$ to compute the $M$-dimensional vector in the metric space of an image, where $\phi$ is the learnable parameters.
-A prototypical representation $\mathbf{c}_{i}$ is defined as the mean of embedded vectors from its support set $\mathcal{S}_{i}$:
-$\mathbf{c}_{i}=\frac{1}{|\mathcal{S}_{i}|}\sum_{\mathbf{x}_{j}\in\mathcal{S}_{i}}f_{\phi}(\mathbf{x}_{j})$ (1)
-After calculating all the $N$ prototypical representations, for an image $\mathbf{x}$ in the query set, we compute its vector $f_{\phi}(\mathbf{x})$ and the distances between it and all the representations $d(f_{\phi}(\mathbf{x}),\mathbf{c}_{i})$, where $d:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}_{0}^{+}$ is a distance function.
-The probability that the query sample $\mathbf{x}$ belongs to class $i$ is calculated by:
-$p(y=i|\mathbf{x})=\mathrm{softmax}_{1\leq i\leq N}(-d(f_{\phi}(\mathbf{x}),\mathbf{c}_{i}))$ (2)
-The image $\mathbf{x}$ is then assigned to the class with maximum probability, or the nearest prototypical representation to it.
-During training, we randomly selected $N_{c}$ classes from the training set $\mathcal{D}_{train}$.
-For each selected class $i$, we randomly choose $N_{s}$ samples as support set $\mathcal{S}_{i}$ and another $N_{q}$ samples as query set $\mathcal{Q}_{i}$.
-We compute the $N_{c}$ prototype centers $\mathbf{c}_{i}$ according to equation 1.
-The prototypical network outputs a probability distribution over the $N_{c}$ classes for each query sample $\mathbf{x}$ in query set.
-The optimization target is to minimize the the negative log-probability $J(\phi)$.
-Pseudo code to compute the loss $J(\phi)$ for a training step is provided in Algorithm 1.
+Our FSD is based on the Prototypical Network (Snell et al., 2017), an effective method for addressing the few-shot question above.
+The Prototypical Network aims to learn a metric space in which samples from the same class are close in distance, while samples from different classes are relatively distant from each other.
+Thus, we can represent each class by a single vector, known as the prototypical representation.
+For each sample in the query set, classification is conducted by simply finding the nearest representation to it.
+In detail, as depicted in Figure 2, FSD adopts a neural network $f_{\phi}:\mathbb{R}^{D}\to\mathbb{R}^{M}$ to compute the $M$-dimensional vector in the metric space of an image, where $\phi$ is the learnable parameters.
+The prototypical representation $\mathbf{c}_{i}$ for class $i$ is defined as the average of the embedded vectors derived from its corresponding support set $\mathcal{S}_{i}$:
+$\mathbf{c}_{i}=\frac{1}{|\mathcal{S}_{i}|}\sum_{\mathbf{x}_{j}\in\mathcal{S}_{i}}f_{\phi}(\mathbf{x}_{j})\text{.}$ (1)
+After gaining all the $N$ prototypical representations and the vector $f_{\phi}(\mathbf{x}_{q})$ of a test image $\mathbf{x}_{q}$ in the query set, we compute the distances between the test vector and all the representations $d(f_{\phi}(\mathbf{x}_{q}),\mathbf{c}_{i})$, using a distance function $d:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}_{0}^{+}$.
+The probability that the query sample $\mathbf{x}_{q}$ belongs to class $i$ is determined by:
+$p(y=i|\mathbf{x}_{q})=\mathrm{Softmax}_{1\leq i\leq N}(-d(f_{\phi}(\mathbf{x}_{q}),\mathbf{c}_{i}))\text{.}$ (2)
+The image $\mathbf{x}_{q}$ is assigned to the class with the maximum probability, or the nearest prototypical representation to it.
+As shown in Figure 3(a), during each training step, $N_{c}$ classes are randomly selected from the training set.
+For each selected class $i$, we randomly choose $N_{s}$ samples from it as the support set $\mathcal{S}_{i}$ and another $N_{q}$ samples as the query set $\mathcal{Q}_{i}$.
+We compute the $N_{c}$ prototypical representations according to Equation 1.
+The Prototypical Network outputs a probability distribution over the $N_{c}$ classes for each sample in the query set.
+The optimization target is to minimize the negative log-probability $J(\phi)$.
+Pseudo code to compute the loss $J(\phi)$ is provided in Algorithm 1.
 Algorithm 1 Prototypical Loss Computation Algorithm
-  Input: Training set $\mathcal{D}=\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\cdots\mathcal{D}_{N}$, where $\mathcal{D}_{k}$ denotes the subset of class $k$, number of training class $N_{C}$, number of support examples per class $N_{S}$, number of query examples per class $N_{Q}$.
-  Output: training loss $J$ of one step.
-  $V\leftarrow\textit{randomly select }N_{c}\textit{ classes from }\{1,2,\ldots,N\}$
+  Input: Training set $\mathcal{D}=\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\cdots\mathcal{D}_{N}$, where $\mathcal{D}_{k}$ denotes the subset of class $k$; number of training class $N_{c}$; number of support examples per class $N_{s}$; number of query examples per class $N_{q}$.
+  Output: training loss $J$.
+  $V\leftarrow\text{randomly select }N_{c}\text{ classes from }\{1,2,\ldots,N\}$
   for $k$ in $V$ do
-     $S_{k}\leftarrow\textit{randomly select }N_{s}\textit{ samples from }\mathcal{D}_{k}$
-     $Q_{k}\leftarrow\textit{randomly select }N_{q}\textit{ samples from }\mathcal{D}_{k}/S_{k}$
+     $S_{k}\leftarrow\text{randomly select }N_{s}\text{ samples from }\mathcal{D}_{k}$
+     $Q_{k}\leftarrow\text{randomly select }N_{q}\text{ samples from }\mathcal{D}_{k}/S_{k}$
      $\mathbf{c}_{k}\leftarrow\frac{1}{N_{s}}\sum_{\mathbf{x}_{i}\in S_{k}}f_{\phi}(\mathbf{x}_{i})$
@@
   for $k$ in $V$ do
-     $J\leftarrow J+\frac{1}{N_{c}N_{q}}\sum_{\mathbf{x}_{j}\in Q_{k}}\log\mathrm{softmax}(-d(f_{\phi}(\mathbf{x}_{j}),\mathbf{c}_{k}))$
+     $J\leftarrow J-\frac{1}{N_{c}N_{q}}\sum_{\mathbf{x}_{j}\in Q_{k}}\log\mathrm{Softmax}(-d(f_{\phi}(\mathbf{x}_{j}),\mathbf{c}_{k}))$
   end for
@@
 3.3 Zero-shot Classification
-We propose a zero-shot detection way to verify the applicability of FSD in real-world settings.
-In the zero-shot task, there is no additional samples from the test classes.
-Instead, we are provided with a meta-data vector for each class in the training set to categorize the unseen images.
-We place considerable emphasis on zero-shot classification, since in practical scenarios, it is unfeasible to obtain test samples from every possible generative model.
+We propose a zero-shot detection approach to verify the applicability of FSD in real-world settings.
+In the zero-shot task, there are no additional samples from the test classes.
+Instead, we are provided with a metadata vector for each class in the training set to categorize the test images.
+We put considerable emphasis on this task, since acquiring samples from all potential generative models can be impractical.
 The zero-shot results can represent the performance of our model in general scenarios.
-In this work, we simply define the meta-data vector as the prototypical representation derived from an extensive collection of samples.
-That is, we randomly select plenty of images from each class in the training set and then compute their prototypical representations following Equation 1.
-Once the nearest representation of a test image is labeled as a synthetic class, it will be regarded as a fake image.
-4 Experiment
-Table 1: Comparison of our FSD and previous studies.
+Table 1: Comparison of our FSD with previous studies.
 We train several classifiers on different subsets and report the accuracy for each unseen test subset.
@@
 Midjourney GLIDE ADM SD VQDM BigGAN Acc(%)
-CNNSpot (Wang et al., 2020) 58.2 57.1 57.0 70.3 56.7 56.6 59.3
-Spec (Zhang et al., 2019b) 56.7 65.4 57.9 72.4 61.7 64.3 63.1
-F3Net (Qian et al., 2020) 55.1 57.8 66.5 73.1 62.1 56.5 61.9
-GramNet (Liu et al., 2020) 58.1 65.3 58.7 72.8 57.8 61.2 62.3
-DIRE (Wang et al., 2023) 65.0 69.1 61.9 73.7 63.4 56.7 65.0
-LARE2 (Luo et al., 2024) 66.4 81.3 66.7 87.3 84.4 74.0 76.7
+Spec (Zhang et al., 2019b) 50.0 64.7 52.8 56.1 56.5 63.0 57.2
+CNNSpot (Wang et al., 2020) 52.8 73.3 55.0 55.9 54.4 66.2 59.6
+F3Net (Qian et al., 2020) 50.1 52.5 66.4 57.0 59.4 50.4 56.0
+GramNet (Liu et al., 2020) 51.3 62.6 53.8 56.8 52.2 57.2 55.7
+DIRE (Wang et al., 2023) 57.9 68.2 57.3 58.2 59.6 50.8 58.7
+LARE2 (Luo et al., 2024) 62.7 80.2 63.5 79.6 76.9 72.0 72.5
 FSD (zero-shot) 75.1 93.9 74.1 88.0 69.1 62.1 77.1
 FSD (10-shot) 80.9 97.1 79.2 88.8 76.2 82.2 84.1
+In this work, we simply define the metadata vector as the prototypical representation derived from an extensive collection of samples.
+That is, we randomly select a large number of images from each class in the training set and then compute the prototypical representations following Equation 1.
+If the nearest representation of a test image is labeled as a synthetic class, the image will be considered fake.
+4 Experiment
 In this section, we first introduce the experimental setups and then provide a comparison between FSD and previous studies.
@@
 The diffusion generators include Midjourney (mid, 2022), Stable Diffusion V1.4 (Rombach et al., 2022), Stable Diffusion V1.5 (Rombach et al., 2022), Wukong (wuk, 2022), ADM (Dhariwal & Nichol, 2021), GLIDE (Nichol et al., 2022) and VQDM (Gu et al., 2022).
-And the only GAN is BigGAN (Brock et al., 2019).
+The only GAN utilized is BigGAN (Brock et al., 2019).
 The real images are first divided into $8$ subsets.
-And each subset is then further partitioned into training part and test part.
-Each generator is paired with a specific subset and uses the category labels of the real images within the subset to produce synthetic images.
-Therefore, GenImage dataset consists of $8$ fake-vs-real subsets corresponding to the $8$ generative models.
-In this study, the synthetic images in different subsets are categorized into distinct classes, whereas the real images across all subsets are classified into a single class.
-We also notice that $3$ diffusion models SD v1.4, SD v1.5 and Wukong share an identical model structure and it is difficult to distinguish among them.
-Thus, we combine their corresponding subsets into a unified subset, named Stable Diffusion (SD).
-Finally, the dataset used in our experiments comprises $7$ classes of images from different sources.
+Each subset is subsequently partitioned into a training part and a test part.
+Each generator is associated with one specific subset and uses the category labels of the real images within the subset to produce synthetic images.
+Consequently, the GenImage dataset consists of $8$ fake-vs-real subsets for analysis.
+In this study, the synthetic images in different subsets are categorized into distinct classes, while real images from all subsets are aggregated into a single class.
+We also observe that $3$ diffusion models, SD v1.4, SD v1.5 and Wukong, share an identical model structure, making it challenging to differentiate among them.
+Thus, we merge their corresponding subsets into a unified subset, named Stable Diffusion (SD).
+Ultimately, the dataset used in our experiments comprises $7$ classes of images from different sources.
 Evaluation metrics.
@@
 Following Wang et al.
-(2023), the input images are first resized to $256\times 256$ and then randomly cropped to $224\times 224$ with a random horizontal flip during training.
+(2023), the input images are first resized to $256\times 256$ and then randomly cropped to $224\times 224$ with random horizontal flipping during training.
 In contrast, only a center crop to $224\times 224$ is performed after resizing the images during testing.
 The distance in the metric space is measured with Squared Euclidean Distance.
-To fit the classification task, we apply different few-shot manners to training and testing.
-In detail, we select one synthetic subset as the test data and utilize the rest $6$ subsets as the training data for one classifier.
-At every training step, $3$ classes are randomly selected from the training set, with $5$ random samples chosen from each class for the support batch and another $5$ samples for the query batch.
-We employ Adam as the optimizer to minimize the cross entropy loss with a base learning rate of $10^{-4}$.
-We also adopt a StepLR scheduler with $\gamma=0.5$ and $step=80000$.
-The training of each classifier is carried out for $200,000$ steps with a batch size of $16$.
-During testing, we conduct a $2$-way classification task by comparing the selected test class against the real class.
-For the zero-shot detection, we randomly select $1024$ samples from each class in the training set to compute the meta-data vectors.
-Our method is implemented by PyTorch library and all the experiments are conducted on a single A100 with 40GB memory.
-Table 2: Result of $10$-shot cross-generator image classification on different subsets.
+To fit the classification task, we employ different few-shot strategies for training and testing.
+Specifically, we select one synthetic subset as the test set and utilize the remaining $6$ subsets to train a classifier.
+At every training step, $3$ classes are randomly selected from the training set, with $5$ samples chosen from each class for the support set and another $5$ samples per class for the query set.
+We employ Adam as the optimizer to minimize the cross-entropy loss with a base learning rate of $10^{-4}$.
+We also adopt a StepLR scheduler with $\gamma=0.5$ and $step\_size=80000$.
+Each classifier is trained for $200,000$ steps, with a batch size of $16$.
+During testing, we conduct a binary classification task by comparing the selected test class against the real class.
+For zero-shot detection, we randomly select $1024$ samples from each training subset to compute the metadata vectors.
+Our method is implemented with the PyTorch library and all the experiments are conducted on a single A100 with 40GB memory.
+Table 2: Results of $10$-shot cross-generator image classification on different subsets.
 Each classifier is trained on the dataset excluding the subset from the first column.
@@
 4.3 Comparison to Existing Methods
-We compare FSD with several state-of-the-art synthetic image detection methods, including CNNSpot (Wang et al., 2020), Spec (Zhang et al., 2019b), F3Net (Qian et al., 2020), GramNet (Liu et al., 2020), DIRE (Wang et al., 2023) and LARE2 (Luo et al., 2024).
+We compare FSD with several state-of-the-art synthetic image detection methods and summarize the results in Table 1.
+The compared group includes CNNSpot (Wang et al., 2020), Spec (Zhang et al., 2019b), F3Net (Qian et al., 2020), GramNet (Liu et al., 2020), DIRE (Wang et al., 2023) and LARE2 (Luo et al., 2024).
 The first four methods are model-agnostic, focusing on identifying synthetic artifacts within fake images.
-And the last two methods leverage the reconstruction error of Stable Diffusion to differentiate between fake and real images.
+The last two methods leverage the reconstruction error of Stable Diffusion to differentiate between fake and real images.
 To conduct a comprehensive comparison, we train $6$ classifiers on different training sets, each of which excludes a specific test subset.
 However, previous studies typically provide several classifiers, and each classifier is trained on one distinct subset and evaluated on the others.
-To minimize the difference between our method and previous researches, we report the average performance of those classifiers over each test subset in previous studies, which can better represent their generalization ability on each class.
-The results are summarized in Table 1.
-As illustrated in Table 1, our FSD attains state-of-the-art performance across $5$ out of the $6$ test classes, surpassing previous state-of-the-art by $+7.4\%$ accuracy in average.
+To minimize the difference between our method and previous works, we report the average performance of $5$ classifiers trained on non-test categories and evaluated on each test subset, which can better represent their generalization ability on each class.
+As illustrated in Table 1, our FSD attains state-of-the-art performance across $5$ out of the $6$ test classes, surpassing previous state-of-the-art by $+11.6\%$ accuracy on average.
 The suboptimal performance of our model on the VQDM class can be attributed to the involvement of image quantization in VQDM, which significantly differs from the other generative models in our experiments.
 Another reason is that LARE2 is specifically designed to detect synthetic images generated by diffusion models, which enables it to achieve unexpectedly high performance on the VQDM class.
-Although our few-shot method may not be directly comparable to other approaches due to the accessing of the few samples in the test class, we still regard this discrepancy as a significant strength of our model.
+Although our few-shot method may not be directly comparable to other approaches due to accessing only a few samples in the test class, we still regard this discrepancy as a significant strength of our model.
 The definite improvement in performance demonstrates FSD’s capability to effectively leverage a limited number of unseen samples.
 Based on Table 1, we find that FSD achieves notable performance in zero-shot scenarios, in which the prototypical representations are calculated from samples in the training classes.
-The result shows that our approach is also suitable for detecting synthetic images even when samples from the same sources are not provided, thus broadening the application scope of FSD.
-We owe this advantage to FSD’s capability to capture not only the intra-class characteristics but also the common features across all classes.
+These results show that our approach is also suitable for detecting synthetic images even when samples from the same sources are not provided, thus broadening the application scope of FSD.
+We attribute this advantage to FSD’s ability to capture not only the intra-class characteristics but also the common features across all classes.
 Another observation is that the discrepancy between zero-shot results and few-shot results remains notable, which indicates the necessity of gathering training data from a wider range of generative models.
@@
 When it comes to detecting the unseen classes, as shown in the diagonal of Table 2, FSD suffers from a slight decline in performance, demonstrating the generalization capability of our model.
-Another finding is that FSD performs nearly perfectly on the VQDM class when this class has been included in the training set.
+Another finding is that FSD performs nearly perfectly on the VQDM class when this class is included in the training set.
 However, the performance significantly declines when this class has been excluded.
-This fact shows that it is still challenging to generalize across unseen generative models, indicating that our approach remains inadequate in capturing the discriminative features between classes.
-We also observe the varying performance of FSD across different classes.
-For instance, images generated by GLIDE, ADM, SD and VQDM are relatively easy to detect by our model when these classes are included in the training set, while those from Midjourney and BigGAN are a little challenging to identify.
-Particularly, images from the GLIDE class can be distinguished with exceptional accuracy, even when not included in the training set.
-These facts indicate the necessity of treating images from different sources as distinct categories, which has not been fully explored in previous researches.
-Moreover, it is essential to collect images from representative generative models for training, as generalization on similar generators is easier than on those that are vastly different.
+This fact shows that it is still challenging to generalize across vastly different unseen generative models, indicating that our approach remains inadequate in capturing the discriminative features between classes.
 (a) Test on Midjourney
@@
 The features are extracted from the final layer of our model.
+We also observe the varying performance of FSD across different classes.
+For instance, images generated by GLIDE, ADM, SD and VQDM are relatively easy to detect by our model when these classes are included in the training set, while those from Midjourney and BigGAN are a little challenging to identify.
+Particularly, images from the GLIDE class can be distinguished with exceptional accuracy, even when not included in the training set.
+These facts indicate the necessity of treating images from different sources as distinct categories, which has not been fully explored in previous research.
+Moreover, it is essential to collect images from representative generative models for training, as generalization on similar generators is easier than on those that are vastly different.
 4.5 Visualization
@@
 We randomly selected $1024$ samples from the test part of each class for visualization.
-And the features are extracted from the final layer of our model.
+The features are extracted from the final layer of our model.
 From Figure 4, we can see that features of images from different generative models exhibit distinct distributions, each of which is marked by a unique color.
@@
 However, the boundary of the Midjourney class represented in red is not as clear as that of the others.
-This can be attributed to its API service, which may utilize multiple generative models for image generation based on our inference.
+This can be attributed to its API service, which may utilize multiple generative models for image generation as inferred from our analysis.
 Although the boundary of the test class in each sub-figure is a little ambiguous, the distribution of this class still remains identifiable, which demonstrates that our method successfully captures the intra-category features of unseen data.
@@
 However, such categorization also introduces the issue of eliminating the significant differences between seen data and unseen data.
-As illustrated in Figure 3(e), The distribution of the VQDM class in orange exhibits a large overlap with that of the real class in gray, due to the great difference in image tokenization between VQDM and other generative models.
-This also indicates that our method still faces the challenge in detecting images from models that are significantly different from those generative models used in training, which will be a main focus of our future work.
+As illustrated in Figure 4(e), the distribution of the VQDM class in orange exhibits a large overlap with that of the real class in gray, due to VQDM’s significant differences in image tokenization compared to other models.
+This also indicates that our method still faces the challenge of detecting images from models that are significantly different from those generative models used in training, which will be the main focus of our future work.
 4.6 Ablation Study
@@
 We choose different numbers of support samples in each class, ranging from $1$ to $200$.
-To present a comprehensive overview of this aspect, we conducted the experiment across all the $6$ classifers described in Section 4.4.
+To present a comprehensive overview of this aspect, we conduct the experiment across all the $6$ classifiers described in Section 4.4.
 We evaluate each classifier on the corresponding test class with varying numbers of support samples.
 To maintain a consistent total number of test samples across different settings, we set a fixed ratio of $1:3$ between the support set and the query set.
-The results are shown in Figure 4, in which each fold line represents the performance of a classifier across varying sizes of support set.
+The results are shown in Figure 5.
+Each line represents the performance of one classifier across varying sizes of support set.
 From Figure 5, we can see that using more support samples results in better performance.
@@
 However, it achieves $81.7\%$ accuracy when using $200$ shots, gaining only an increase of $+2.5\%$ over the $10$-shots scenario.
-As it is easy to collect $10$ samples from online generative models and the computation consumption is very little, We ultimately decide to use the $10$-shots results of FSD for comparison.
+As it is easy to collect $10$ samples from online generative models and the computational cost is very low, we ultimately decide to use the $10$-shots results of FSD for comparison.
 1111333355551010101025252525505050501001001001002002002002005050505060606060707070708080808090909090100100100100Number of shots per classAccuracy (%)MidjourneyGLIDEADMSDVQDMBigGAN
-Figure 4: Influences of the number of shots. Each classifier is trained on those subsets excluding the test one. And $10$-shots is the most cost-effective setting across all the categories.
+Figure 5: Influence of the number of shots. Each classifier is trained on those subsets excluding the test one. The $10$-shot setting is the most cost-effective setting across all categories.
+Impact of the multi-class classification.
+Different from previous studies that treat the fake images as a single class and train a binary classifier to detect them, we categorize the images from different generators as distinct classes.
+We argue that distinguishing images from different generators can also benefit the detection performance.
+Thus, we train another binary classifier and compare it with our method on the ADM class, which is difficult to distinguish according to our experiments.
+Both classifiers utilize the ResNet-$50$ as the backbone and both of them are trained on the dataset excluding the ADM class.
+We also apply the t-SNE visualization and randomly select $1024$ samples for visualizing.
+Features of the binary classifier are extracted from the output layer of the ResNet-$50$, with dimensions of $2048$.
+The visualization results are shown in Figure 6.
+Figure 6(a) shows that our FSD successfully distinguishes those images across different classes which have been used in training.
+Images from the unseen ADM class colored in purple also cluster together in the visualization picture.
+In contrast, the traditional binary classifier fails to distinguish images from Midjourney and Stable Diffusion, as shown in Figure 6(b).
+The unseen samples in purple are quite dispersed, failing to form a compact cluster.
+Thus, distinguishing them from the real ones can be extremely difficult.
+This fact indicates that our method possesses a certain ability to extract intra-category commonality from unseen classes, thereby enabling an accurate differentiation among various classes.
+Moreover, we find that it is difficult for a classifier to distinguish samples from Stable Diffusion V1.4 and Stable Diffusion V1.5.
+One plausible explanation could be the shared usage of an identical backbone and VAE, which makes their products highly similar.
+This implies that we should carefully select the representative generative models to build the training set, which we leave for future work.
 (a) FSD
 (b) Binary Classifier
-Figure 5: Visualization of network features for our FSD (a) and a binary classifier (b) on ADM class.
+Figure 6: Visualization of network features for our FSD (a) and a binary classifier (b) on the ADM class.
 In the feature space of FSD, images from the unseen class in purple are tightly clustered together.
-However, they are much more dispersed in the feature space of the binary classifier.
-Impact of the multi-class classification.
-Different from previous studies which treat the fake images as a single class and train an binary classifier to detect them, we category the images from different generators as distinct classes.
-We argue that distinguishing images from different generators can also benefit the detection performance.
-Thus, we train another binary classifier and compare it with our method on the ADM class, which is difficult to detect according to our experiments.
-Both classifiers utilize the ResNet-$50$ as backbone and are trained on the dataset excluding the ADM class.
-We also apply the t-SNE visualization and random select $1024$ samples for visualizing.
-Features of the binary classifier are extracted from the output layer of the ResNet-$50$, with a dimension of $2048$.
-The visualization results are shown in Figure 5.
-Figure 5(a) shows that our FSD successfully distinguishes those images across different classes which has been used in training.
-Images from the unseen ADM class dyed in purple also cluster together in the visualization picture.
-In contrast, the traditional binary classifier fails to distinguish images from Midjourney and Stable Diffsion, as shown in Figure 5(b).
-And the unseen samples in purple are quite dispersed, failing to form a compact cluster.
-Thus, distinguishing them from the real ones can be extremely difficult.
-This fact indicates that our method is able to extract intra-category common features from unseen classes, thereby enabling an accurate differentiation among various classes.
-Moreover, we find that it is difficult for a classifier to distinguish samples from Stable Diffusion V1.4 and Stable Diffusion V1.5.
-One plausible explanation could be the shared usage of an identical backbone and VAE, which makes their products significantly similar.
-This implies that we should carefully select the representative generative models to build the training set, which We leave for future work.
+However, they are significantly more dispersed in the feature space of the binary classifier.
 5 Conclusion and Limitations
-The key point of detecting fake images in real-world scenarios is to make full use of the limited images from unseen generative models.
-In this paper, we propose Few-Shot Detector (FSD), a novel AI-generated image detector based on few-shot learning.
-Experiments show that FSD achieves state-of-the-art performance in synthetic image detection with only $10$ samples from the test generator, demonstrating the strong generalization ability of our model.
-We visualize the feature space to highlight the advantage of distinguishing images from different sources, which has not been widely adopted among previous studies.
+The key point of detecting fake images in real-world scenarios is to make full use of the limited number of images from unseen generative models.
+In this paper, we propose the Few-Shot Detector (FSD), a novel AI-generated image detector based on few-shot learning.
+Experiments show that FSD achieves state-of-the-art performance in synthetic image detection with only $10$ samples from the test generators, demonstrating the strong generalization ability of our model.
+We visualize the feature space to highlight the advantage of categorizing synthetic images by their sources, an approach that has not been widely adopted in previous studies.
 However, our method still suffers from a few limitations.
-Since it is not feasible to obtain few-shot samples from all generative models, our collection of synthetic images is confined to part of them, which may lead to a decrease in detection performance.
+Since it is not feasible to obtain few-shot samples from all generative models, our collection of synthetic images is confined to part of these models, which may lead to a decrease in detection performance.
 Another weakness is that our method does not perform satisfactorily when detecting images from a vastly different generative model.
@@
 Acknowledgments
-This work is supported by the National Science and Technology Major Project (No. 2022ZD0116314) and the National Science Foundation of China (No. 62106249).
+This work is supported by the National Science and Technology Major Project (2022ZD0116314), the National Natural Science Foundation of China (62106249, 62476070, 6243000159, 62102416), Shenzhen Science and Technology Program (JCYJ20241202123503005, GXWD20231128103232001), Department of Science and Technology of Guangdong (2024A1515011540), and the Natural Science Foundation of Jiangsu Province under Grant BK20243051.
 Impact Statement
-AI-generated image detection has been a subject of extensive researches and is regarded as an important technology to reduce cheating, misinformation and harm on internet platforms and social media.
-Our work aims at enhancing fake image detection performance and achieving better generalization with a few samples, rather than increasing synthetic data.
-All the images used in our experiment are from open-source dataset.
-We believe our work can contribute to the development of deepfake detecting applications in many real-world scenarios.
+AI-generated image detection has been a subject of extensive research.
+It is regarded as an important technology to reduce cheating, misinformation and harm on Internet platforms and social media.
+This work focuses on enhancing the performance of fake image detection rather than increasing synthetic data.
+All the images used in our experiments are from open-source datasets.
+We believe that our work can significantly contribute to the development of deepfake-detection applications in many real-world scenarios.